OpenStudy (snowcrystal):

PLEASE HELP ME IM ON A BIT OF A TIME CRUNCH I'LL FAN AND MEDAL QUESTION BELOW

OpenStudy (anonymous):

yah!

OpenStudy (snowcrystal):

Simplify. \[\frac{ 2 }{ \sqrt{5} }\]

rvc (rvc):

rationalize

OpenStudy (anonymous):

is that the complete problem?

OpenStudy (snowcrystal):

\[A.) 2\sqrt{5}\] \[B.) \sqrt{10}\] \[C.) \frac{ \sqrt{10} }{ 5 }\] \[D.) \frac{ 2\sqrt{5} }{ 5 }\]

OpenStudy (anonymous):

Its D

OpenStudy (yanasidlinskiy):

its d

OpenStudy (snowcrystal):

Yes thats the whole problem and I dont know how to do that and Thaks @sylesh98 and @YanaSidlinskiy

OpenStudy (anonymous):

You want to first get rid of the square root portion. The way we get rid of this is by 2√5/5

OpenStudy (anonymous):

by multiplying 5^1/2 in numerator and denoominator it will give option D

OpenStudy (yanasidlinskiy):

np

rvc (rvc):

i said rationalize the denominator @SnowCrystal

OpenStudy (anonymous):

Yeah what they said. But for future purposes you just rationalize

rvc (rvc):

r u there plz sovle as im saying u can do many problems further

rvc (rvc):

@SnowCrystal

rvc (rvc):

simply:\[63\div \sqrt{3}\]

rvc (rvc):

@SnowCrystal y not answering? not interested plz say i'll not waste time

OpenStudy (snowcrystal):

I'm sorry I had to help my mom

rvc (rvc):

its ok so ans now

rvc (rvc):

hey @SnowCrystal

OpenStudy (snowcrystal):

yes? um what do u mean answer what?

rvc (rvc):

the above q @SnowCrystal

OpenStudy (snowcrystal):

no i am im just trying to rush through and catch up with my class im 12 assignments behind and have to get done by monday

OpenStudy (snowcrystal):

im sorry but i cant at the current moment

rvc (rvc):

Ok as ur wish ALL THE BEST!!!!!!!!!!!!

OpenStudy (snowcrystal):

thanks